A bullet of mass 20 g is fired from a rifle with a velocity of 800 m/s. After passing through a mud wall 100 cm thick, velocity drops to 100 m/s. What is the average resistance force of the wall :(

Grade:12th pass

5 Answers

Mudit
7 years ago
mass=0.02kg u=800m/s v=100m/s
length of wall = dist. covered;s=100cm=1m
v2-u2=2as
(10000-640000)2*1=a
a=-31500m/s2
Avg. resisteive force=m*a
                             =0.02*31500=15750000N
Vishal bhati
5 years ago
Mass=0.20kg
V=800m/s
U=100m/s
Length of the wall=dist. Covered=100cm=1m
V^2-u^2=2as
(100)^2-(800)^2=2×1×a
a=-63000/2
a=-31500
Avg.resistance =m×a
=>0.20kg×31500
F=6300N
Asfand
5 years ago
A bullet of mass 20grams is fired with a muzzle velocity q00ms.find the recoil of gun if its mass is 5 kg
Devang
5 years ago
Mass of bullet = 20g = 0.02 kg
Initial velocity of bullet U = 800m/s
final velocity of bullet V = 100 m/s
Distance covered by bullet = thickness of wall = s=100cm = 1m
using formula V^2 - U^2 = 2as
            (10000-640000)=2*a*1
then a= - 31500 m/s2.
resistance force = m*a
                              = 0.02*31500
                              = 630N (opposite direction of velocity of bullet)
 
 
V
Akshaya
3 years ago
6300 Newton is the answer for this question just substitute the value in third equation of motion v2-u2=2as put value of a= 1m/s2